OpenStudy (anonymous):

In a batch of 960 calculators, 8 were found to be defective. What is the probability that a calculator chosen at random will be defective? Write the probability as a percent. Round to the nearest tenth of a percent if necessary.

OpenStudy (anonymous):

1:120 or like 8.3%

OpenStudy (anonymous):

this is simply 8/960 = 1/120 or 1/120 * 100 = 8.33 %

OpenStudy (kira_yamato):

P(def) = 8/960 (Simplify this)